OpenStudy (babynini):

a) sketch curve b) find a rectangular coordinate equation for the curve by eliminating the perimeter x=sec(t), y=tan(t), 0

OpenStudy (babynini):

I simplified x into x= 1/cos(t) and y = sin(t)/cos(t) then y = [sin(t)][1/cos(t)]

OpenStudy (babynini):

then plugged x into y and got y=sint(x)

OpenStudy (babynini):

eem..I got confused on that second part there because there's lots of ='s xD yeah the first part is easy

OpenStudy (loser66):

ok, line by line y^2 = tan^2 t , ok?

OpenStudy (babynini):

right

OpenStudy (babynini):

yup

OpenStudy (babynini):

yeah.aah I see.

OpenStudy (loser66):

ok, last line is obvious. right?

OpenStudy (babynini):

Yep!!

OpenStudy (babynini):

well the sec^2t= x^2 makes sense

OpenStudy (babynini):

how you got to x^2-1=y^2 I don't know.

OpenStudy (loser66):

how cannot? y^2 +1 = x^2 just -1 both sides

OpenStudy (babynini):

Oooh sorry, i had written it wrong on my paper. k.

OpenStudy (babynini):

it could also be written x^2-y^2=1 right?

OpenStudy (loser66):

sure

OpenStudy (babynini):

then we just plug in numbers for x and find y.

OpenStudy (babynini):

Kk's thanks! :)

OpenStudy (loser66):

hihihi... tip for a) x^2 -y^2 =1 is a hyperbola with the center at origin

OpenStudy (babynini):

Oo that is helpful.

OpenStudy (babynini):

Hold up, how do I find the perimeter?

OpenStudy (babynini):

like they give me 0 </= t </= pi/2 but I have to do that for x and y